# Multi-Level Monte Carlo Training of Neural Operators

### v1

@@ Abstract
Operator learning is a rapidly growing field that aims to approximate nonlinear operators related to partial differential equations (PDEs) using neural operators. These rely on discretization of input and output functions and are, usually, expensive to train for large-scale problems at high-resolution. Motivated by this, we present a Multi-Level Monte Carlo (MLMC) approach to train neural operators by leveraging a hierarchy of resolutions of function dicretization. Our framework relies on using gradient corrections from fewer samples of fine-resolution data to decrease the computational cost of training while maintaining a high level accuracy. The proposed MLMC training procedure can be applied to any architecture accepting multi-resolution data. Our numerical experiments on a range of state-of-the-art models and test-cases demonstrate improved computational efficiency compared to traditional single-resolution training approaches, and highlight the existence of a Pareto curve between accuracy and computational time, related to the number of samples per resolution.
@@ 1 Introduction
Neural operators have been recently introduced as a mathematical tool to learn infinite-dimensional operators, mapping functions to functions, and have been applied in many scientific areas to approximate the solution operator of partial differential equations (PDEs) or solve inverse problems (Li et al., 2021b; Lu et al., 2021; Kovachki et al., 2023; Boullé and Townsend, 2024; Kovachki et al., 2024). Several neural operator architectures have been proposed including Fourier neural operators (FNOs) (Li et al., 2021b), deep operator networks (DeepONets) (Lu et al., 2021), graph neural operators (Pfaff et al., 2020; Anandkumar et al., 2020) and transformers (Cao, 2021; Hao et al., 2023). These networks exploit (or enforce) different properties of operators such as regularity of solutions and geometry of the underlying spatial domain on which functions are defined, and are supported by various approximation theory results (Kovachki et al., 2021, 2024; Lanthaler et al., 2022).
@@
The simulation of physical systems often relies on mesh-based representations, where the choice of mesh resolution presents a fundamental trade-off between computational cost and accuracy (this appears explicitly in, e.g., finite element error estimations (Brenner and Scott, 2008)). Approximated solutions discretized on fine meshes are more accurate, but they significantly increase computational overhead during both training and inference. Although traditional deep learning approaches typically operate at a fixed resolution, limiting their ability to efficiently learn across multiple scales, neural operators are generally designed to process functions defined on arbitrary meshes by learning a representation with respect to a given basis (Li et al., 2021b; Bouziani and Boullé, 2024) and interpolating/projecting across different resolutions. One of the main challenges in scaling neural operators to large-scale three-dimensional problems is the computational cost of the training phase, due to the fine gradient estimations needed to update the parameters of the neural network with a gradient-based optimization algorithm. Little work has been done in the literature on efficient training of neural operators, with the exception of a few works that proposed multi-scale architectures (Li et al., 2020; He et al., 2024; Bouziani and Boullé, 2024).
@@
This work addresses the problem of the high training cost of neural operators on high-resolution data by proposing a flexible gradient estimation algorithm based on Multi-Level Monte Carlo (MLMC) methods (Giles, 2008; Giles and Waterhouse, 2009; Cliffe et al., 2011; Giles, 2015) cast in the framework of mini-batch stochastic optimization. Our framework allows us to estimate the expectation of the gradients of the loss function on the finest resolution with respect to the neural network parameters using a multi-resolution approach. This quantity is expressed as a telescopic sum of gradients expectations across mesh levels. The aim is to drastically decrease the training computational time by capturing the most of the dynamics, i.e., the variance, on the coarse grid, hence decreasing the number of gradient computations needed on the finest grid. We proved that the main assumptions of the MLMC theory are satisfied, under suitable regularity properties, by neural operators thus providing a robust mathematical justification in its application to operator learning. Moreover, our algorithm introduces a notion of “pareto optimality” in the training process, where one can adjust the number of resolution levels (and hence the number of samples across levels) to achieve a trade-off between training time computational cost and accuracy.
Related Works.
@@
Our work translates the classic MLMC methods for PDEs (Giles, 2008; Giles and Waterhouse, 2009; Cliffe et al., 2011; Giles, 2015) to the context of neural operators, where the regularity properties of the PDE solution theoretically justify the approximation of the gradients. MLMC methods also inspired the development of several related works in machine learning. In (Li et al., 2021a) an active learning strategy is employed to dynamically query an acquisition function to process new examples at different resolutions. Lye et al. (2021) propose a multi-level approach to learn input parameters to observable maps and apply it to uncertainty quantification. Moreover, Jin et al. (2023) study a multi-level operator learning algorithm to obtain theoretical sample complexity estimates. More recently, Gal et al. (2025); Zakariaei et al. (2025) introduce a multi-scale stochastic gradient descent approach to train convolutional and graph neural networks at different resolutions.
Main contributions.
The key advantages and novelties of this work are summarized below:
  • Introduction of a rigorous mathematical framework supporting the adoption of MLMC training of neural operators at different resolutions.
  • Exploitation of multi-resolution data for the acceleration of training with a single high-resolution dataset, by coarsening, or when access to high-resolution data is limited or scarce.
  • Implementation of data batching engine that enables effective training with multi-resolution dataset and allows different sampling strategies.
  • Reduction of training time by up to 60% compared to single-resolution training across multiple state-of-the-art neural operators architectures and test problems.
@@ 2 Multi-Level Monte Carlo method
Our approach extends the MLMC framework to neural networks training. In particular, we aim at accelerating the training of neural operators by using a multi-resolution approch, i.e., the MLMC paradigm. MLMC methods stand on the idea that most of the uncertainty can be captured on the coarse grids and so the number of samples needed on the finest grid is greatly reduced.
@@ 2.1 Mathematical Framework
We follow the neural operator framework introduced by Kovachki et al. (2023) and consider two Hilbert spaces of functions $\mathcal{A}$ and $\mathcal{U}$, respectively, defined on bounded domains $\Omega_{\mathcal{A}}\subset\mathbb{R}^{d}$ and $\Omega_{\mathcal{U}}\subset\mathbb{R}^{d}$. We aim to approximate a (nonlinear) operator $\mathcal{G}^{\dagger}:\mathcal{A}\to\mathcal{U}$ mapping between these spaces using a neural operator $\mathcal{G}_{\theta}:\mathcal{A}\to\mathcal{U}$, parametrized by a finite number of parameters $\theta\in\mathbb{R}^{p}$. We train the neural operator using pairs of observations $\{a^{(j)},u^{(j)}\}_{j=1}^{N}$, where $a^{(j)}\sim\mu$ are sampled i.i.d. from a probability measure $\mu$ supported on $\mathcal{A}$ (usually a Gaussian measure), and aim to control the $L^{2}_{\mu}(\mathcal{A};\mathcal{U})$ Bochner norm as
[…]
Here, we assume that both $\mathcal{G}^{\dagger}$ and $\mathcal{G}_{\theta}$ are $\mu$-measurable, belong to $L^{2}_{\mu}(\mathcal{A};\mathcal{U})$, and wish to solve the following optimization problem:
@@
$\theta^{k+1}=\theta^{k}-\alpha\nabla_{\theta}(\mathbb{E}_{a\sim\mu}\|\mathcal{G}^{\dagger}(a)-\mathcal{G}_{\theta}(a)\|_{\mathcal{U}}^{2}),$ (1)
where we move the gradient inside the expectation using the dominated convergence theorem (see. Eq. 4 in Appendix A), and denote by $X_{\theta}:\mathcal{A}\to\mathbb{R}$ the operator defined as $X_{\theta}(a)=\|\mathcal{G}^{\dagger}(a)-\mathcal{G}_{\theta}(a)\|_{\mathcal{U}}^{2}$.
Function space hierarchy.
For $1\leq i\leq m$, we define $\mathcal{A}_{i}$ to be a finite element space of functions defined on a mesh $\Omega_{\mathcal{A}}^{i}$ of $\Omega_{\mathcal{A}}$, with polynomial degree $k\geq d/2$, and mesh diameter $h_{i}=2^{-i}h_{0}>0$ such that $\dim(\mathcal{A}_{1})\leq\cdots\leq\dim(\mathcal{A}_{m})$ (note that one could use another refinement factor). For simplicity, we assume that $\mathcal{A}_{i}$ are subspaces of $\mathcal{A}$ so that the injection operator $\mathcal{A}_{i}\to\mathcal{A}$ is trivial, and define the Galerkin projection operator $\mathcal{P}_{i}:\mathcal{A}\to\mathcal{A}_{i}$ and $X_{\theta}^{i}=X_{\theta}\circ\mathcal{P}_{i}$. Note that the neural operator $\mathcal{G}_{\theta}$ is discretization invariant following the framework introduced in (Kovachki et al., 2023), and that computing $X_{\theta}^{1}\circ\mathcal{P}_{1}$ (and its gradient with respect to $\theta$) is inherently cheaper than computing $X_{\theta}^{m}\circ\mathcal{P}_{m}$. Hence, the projection step to the latent dimension of the neural operator is computationally cheaper. Note that this applies for any neural operator architecture written in the encoder-processer-decoder framework (Bouziani and Boullé, 2024; Kovachki et al., 2024). FNOs (Li et al., 2021b) could also be written in this framework by considering a spectral discretization of the function space, and the Fourier transform as the encoder.
Multi-Level Monte Carlo gradient estimation.
The goal of the MLMC (Giles, 2015) gradient estimation is to approximate the expectation of the gradient of the loss function with respect to the neural network parameters $\theta$ using a telescopic sum of differences across mesh levels. The main idea is to use a coarse mesh to capture most of the variance in the gradient, and then use finer meshes to correct for the remaining variance. This allows us to reduce the number of fine-mesh samples needed, thus speeding up the training process. We express the fine-mesh gradient expectation as a telescopic sum of differences across mesh levels and decompose the expectation of the gradient in Eq. 4 as follows:
@@
$\approx\frac{1}{N_{1}}\sum_{j=1}^{N_{1}}[\nabla_{\theta}X_{\theta}^{1}(a_{1,j})]+\sum_{i=2}^{m}\left(\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}[\nabla_{\theta}X_{\theta}^{i}(a_{i,j})-\nabla_{\theta}X_{\theta}^{i-1}(a_{i,j})]\right),$
@@
where $\{N_{i}\}_{1\leq i\leq m}$ denotes the sample sizes at each level. Note that we do not consider discretization of the output space $\mathcal{U}$ for simplicity, but a similar argument would also work, see, e.g., (Bouziani and Boullé, 2024). The following proposition relies on a standard MLMC result (Giles, 2008; Giles and Waterhouse, 2009) and provides a mean squared error bound for the MLMC estimator in Eq. 2. The proof is deferred to Appendix A.
Proposition 2.1.
@@
Assume that $\mu$ is compactly supported on $H^{k}(\Omega_{\mathcal{A}})$, and that the cost of each sample in Eq. 2 scales linearly with respect to the number of degrees of freedom in the finite element space as $C_{i}=\mathcal{O}(k^{d}2^{di})$ for $1\leq i\leq m$. For any $\epsilon>0$, under enough regularity on $\mathcal{G}^{\dagger}$ and $\mathcal{G}_{\theta}$, there exist a number of levels $m>0$ and an optimal allocation of samples $N_{i}\propto 2^{-(2k+d)i/2}$ per level such that the MLMC estimator given by Eq. 2 has a mean squared error bounded by $\epsilon^{2}$.
@@
Following Proposition 2.1, we find that one can reduce the number of fine-mesh samples to accurately estimate the gradient of the loss function by adapting the number of samples at each level depending on the mesh resolution. Note that the number of samples per resolution is related to the sample cost assumption, which depends on the type of neural operator architecture. The analysis can also be generalized to non-geometric hierarchy of meshes (Haji-Ali et al., 2016).
@@
(Multi-Level Quasi-Monte Carlo) The computational cost of MLMC can be reduced by employing a Quasi-Monte Carlo (QMC) approach, where a nested set of samples across mesh resolutions is chosen instead of independent realizations at each level (Kuo et al., 2012, 2015). This increases the variance of the estimator but results in a lower cost complexity (Giles, 2015).

### v2

@@ Abstract
Operator learning is a rapidly growing field that aims to approximate nonlinear operators related to partial differential equations (PDEs) using neural operators. These rely on discretization of input and output functions and are, usually, expensive to train for large-scale problems at high-resolution. Motivated by this, we present a Multi-Level Monte Carlo (MLMC) approach to train neural operators by leveraging a hierarchy of resolutions of function discretization. Our framework relies on using gradient corrections from fewer samples of fine-resolution data to decrease the computational cost of training while maintaining a high level accuracy. The proposed MLMC training procedure can be applied to any architecture accepting multi-resolution data. Our numerical experiments on a range of state-of-the-art models and test-cases demonstrate improved computational efficiency compared to traditional single-resolution training approaches, and highlight the existence of a Pareto curve between accuracy and computational time, related to the number of samples per resolution.
keywords:
Operator Learning , Neural Operator , Multi-Level Monte Carlo , Multi-Resolution Data
journal: CMAME
[…]
[label1]organization=Department of Applied Mathematics and Theoretical Physics, University of Cambridge,city=Cambridge, postcode=CB3 0WA, country=UK
[…]
[label2]organization=Department of Mechanical Engineering, University of Washington,city=Seattle, postcode=WA 98195, country=USA
[…]
[label3]organization=Department of Computer Science and Technology, University of Cambridge,city=Cambridge, postcode=CB3 0WA, country=UK
[…]
[label4]organization=Department of Mathematics, Imperial College London,city=London, postcode=SW7 2AZ, country=UK
@@ 1 Introduction
Neural operators have been recently introduced as a mathematical tool to learn infinite-dimensional operators, mapping functions to functions, and have been applied in many scientific areas to approximate the solution operator of partial differential equations (PDEs) or solve inverse problems [32, 34, 25, 3, 23]. Several neural operator architectures have been proposed including Fourier neural operators (FNOs) [32], deep operator networks (DeepONets) [34], graph neural operators [36, 1] and transformers [7, 17]. These networks exploit (or enforce) different properties of operators such as regularity of solutions and geometry of the underlying spatial domain on which functions are defined, and are supported by various approximation theory results [24, 23, 28].
@@
The simulation of physical systems often relies on mesh-based representations, where the choice of mesh resolution presents a fundamental trade-off between computational cost and accuracy (this appears explicitly in, e.g., finite element error estimations [6]). Approximated solutions discretized on fine meshes are more accurate, but they significantly increase computational overhead during both training and inference. Although traditional deep learning approaches typically operate at a fixed resolution, limiting their ability to efficiently learn across multiple scales, neural operators are generally designed to process functions defined on arbitrary meshes by learning a representation with respect to a given basis [32, 4] and interpolating/projecting across different resolutions. One of the main challenges in scaling neural operators to large-scale three-dimensional problems is the computational cost of the training phase, due to the fine gradient estimations needed to update the parameters of the neural network with a gradient-based optimization algorithm. Little work has been done in the literature on efficient training of neural operators, with the exception of a few works that proposed multi-scale architectures [31, 18, 4] or adaptive subsampling [29].
[…]
Figure 1: Schematic diagram of the Multi-Level Monte Carlo (MLMC) approach for training neural operators. The input and output training data is downsampled from fine to coarse resolutions. The MLMC gradient estimation in Eq. 3 combines samples from different levels to reduce the overall training cost while maintaining accuracy.
@@
This work addresses the problem of the high training cost of neural operators on high-resolution data by proposing a flexible gradient estimation algorithm based on Multi-Level Monte Carlo (MLMC) methods [14, 13, 8, 15] cast in the framework of mini-batch stochastic optimization. Our framework allows us to estimate the expectation of the gradients of the loss function on the finest resolution with respect to the neural network parameters using a multi-resolution approach. This quantity is expressed as a telescopic sum of gradients expectations across mesh levels. The aim is to drastically decrease the training computational time by capturing most of the dynamics, i.e., the variance, on the coarse grid, hence decreasing the number of gradient computations needed on the finest grid. We proved that the main assumptions of the MLMC theory are satisfied, under suitable regularity properties, by neural operators thus providing a robust mathematical justification in its application to operator learning. Moreover, our algorithm introduces a notion of “Pareto optimality” in the training process, where one can adjust the number of resolution levels (and hence the number of samples across levels) to achieve a trade-off between training time computational cost and accuracy.
@@
Our work translates the classic MLMC methods for PDEs [14, 13, 8, 15] to the context of neural operators, where the regularity properties of the PDE solution theoretically justify the approximation of the gradients. MLMC methods also inspired the development of several related works in machine learning. In [30] an active learning strategy is employed to dynamically query an acquisition function to process new examples at different resolutions. [35] proposes a multi-level approach to learn input parameters to observable maps and apply it to uncertainty quantification. Moreover, [20] studies a multi-level operator learning algorithm to obtain theoretical sample complexity estimates. More recently, [11, 44] introduce a multi-scale stochastic gradient descent approach to train convolutional and graph neural networks at different resolutions. In particular, [44] focuses on multi-scale training for standard CNNs applied to image processing tasks, while our work targets neural operators for challenging PDE problems using a variety of neural operator architectures.
While MLMC is a classical variance reduction technique, its extension to neural operator training is challenging. We develop a rigorous mathematical framework in Section 2.1 that adapts MLMC theory to the operator learning setting, proving that neural operators satisfy key MLMC assumptions (cf. Proposition 2.2). Building on this theory, we embed the MLMC estimator into mini-batch stochastic optimization by devising novel multi-resolution sampling and batching strategies in Section 2.2. This includes introducing sample allocation formulas and sub-sampling schemes, which together enable efficient and consistent multi-resolution gradient estimation within each mini-batch. Implementing these ideas required significant engineering effort; indeed, we designed a cache-aware, memory-efficient multi-resolution batcher that manages data across mesh levels and minimizes costly memory transfers. Finally, we demonstrate the generality and effectiveness of our approach through extensive benchmarks in Section 3 on diverse state-of-the-art neural operator architectures (FNO, a message-passing PDE solver, and GINOT) and multiple challenging datasets (including 2D Darcy flow, time-dependent Navier–Stokes, and a 3D jet-engine bracket problem). These experiments show that our MLMC-based training scheme consistently accelerates convergence with minimal loss in accuracy.
@@ 2 Multi-Level Monte Carlo method
Our approach extends the MLMC framework to neural networks training. In particular, we aim at accelerating the training of neural operators by using the MLMC paradigm. MLMC methods stand on the idea that most of the uncertainty can be captured on the coarse grids and so the number of samples needed on the finest grid is greatly reduced.
@@ 2.1 Mathematical Framework
We follow the neural operator framework introduced by [25] and consider an input function space $\mathcal{A}=L^{2}(\Omega_{A})$ and an output Hilbert space of functions $\mathcal{U}$, respectively defined on bounded domains $\Omega_{\mathcal{A}}\subset\mathbb{R}^{d}$ and $\Omega_{\mathcal{U}}\subset\mathbb{R}^{d}$. Here, $\mathcal{U}$ is typically a (high-order) Sobolev space $H^{k}(\Omega_{\mathcal{U}})$ for some $k\geq 0$. We aim to approximate a (nonlinear) operator $\mathcal{G}^{\dagger}:\mathcal{A}\to\mathcal{U}$ mapping between these spaces using a neural operator $\mathcal{G}_{\theta}:\mathcal{A}\to\mathcal{U}$, parametrized by a finite number of parameters $\theta\in\mathbb{R}^{p}$. We train the neural operator using pairs of observations $\{a^{(j)},u^{(j)}\}_{j=1}^{N}$, where $a^{(j)}\sim\mu$ are sampled i.i.d. from a probability measure $\mu$ supported on $\mathcal{A}$, and compactly supported on $H^{k}(\Omega_{\mathcal{A}})$ for some $k\geq 0$. We aim to control the $L^{2}_{\mu}(\mathcal{A};\mathcal{U})$ Bochner norm as
$\|\mathcal{G}^{\dagger}-\mathcal{G}_{\theta}\|_{L^{2}_{\mu}(\mathcal{A};\mathcal{U})}^{2}=\mathbb{E}_{a\sim\mu}\|\mathcal{G}^{\dagger}(a)-\mathcal{G}_{\theta}(a)\|_{\mathcal{U}}^{2}=\int_{\mathcal{A}}\|\mathcal{G}^{\dagger}(a)-\mathcal{G}_{\theta}(a)\|_{\mathcal{U}}^{2}\,\textup{d}\mu(a).$
In this section, we assume that the operators $\mathcal{G}^{\dagger}$ and $\mathcal{G}_{\theta}$ satisfies the regularity conditions given by 2.1. While these are strong assumptions, they are standard in the operator learning approximation theory literature [24, 4] and are necessary to control the image of the source term distribution through the nonlinear operator $\mathcal{G}^{\dagger}$.
Assumption 2.1 (Regularity of $\mathcal{G}^{\dagger}$ and $\mathcal{G}_{\theta}$).
The operators $\mathcal{G}^{\dagger}$ and $\mathcal{G}_{\theta}$ are $\mu$-measurable, and belong to $L^{2}_{\mu}(\mathcal{A};\mathcal{U})$. Moreover, $\nabla_{\theta_{i}}\mathcal{G}_{\theta}\in L^{2}_{\mu}(\mathcal{A};\mathcal{U})$, for all $1\leq i\leq p$. We assume that the operators $\mathcal{G}^{\dagger}$, $\mathcal{G}_{\theta}$ are Lipschitz continuous from $\mathcal{A}\to\mathcal{U}$. Moreover, we assume that each component of $\nabla_{\theta}\mathcal{G}_{\theta}$ (with respect to $1\leq s\leq p$) is Lipschitz continuous from $\mathcal{A}\to\mathcal{U}$, with constant $L_{\theta}$.
We wish to solve the following optimization problem:
@@
$\theta^{k+1}$ $=\theta^{k}-\alpha\nabla_{\theta}(\mathbb{E}_{a\sim\mu}[\|\mathcal{G}^{\dagger}(a)-\mathcal{G}_{\theta}(a)\|_{\mathcal{U}}^{2}])$ (1)
$=\theta^{k}-\alpha\mathbb{E}_{a\sim\mu}[\nabla_{\theta}\|\mathcal{G}^{\dagger}(a)-\mathcal{G}_{\theta}(a)\|_{\mathcal{U}}^{2}],$
where the second equality comes from the dominated convergence theorem. Hence, under 2.1, $\nabla_{\theta_{i}}\mathcal{G}_{\theta}\in L^{2}_{\mu}(\mathcal{A};\mathcal{U})$, for all $1\leq i\leq p$, and if $a\sim\mu$, Cauchy–Schwarz inequality yields
Following Cauchy–Schwarz inequality on the expectation of the product of random variables, we can bound the expectation of the gradient as
$\nabla_{\theta}(\mathbb{E}_{a\sim\mu}\|\mathcal{G}^{\dagger}(a)-\mathcal{G}_{\theta}(a)\|_{\mathcal{U}}^{2})=\mathbb{E}_{a\sim\mu}[\nabla_{\theta}\|\mathcal{G}^{\dagger}(a)-\mathcal{G}_{\theta}(a)\|_{\mathcal{U}}^{2}].$ (2)
Function space hierarchy
For $1\leq i\leq m$, we define $\mathcal{A}_{i}$ to be a finite element space of functions defined on a mesh $\Omega_{\mathcal{A}}^{i}$ of $\Omega_{\mathcal{A}}$, with polynomial degree $k\geq d/2$, and mesh diameter $h_{i}=2^{-i}h_{0}>0$ such that $\dim(\mathcal{A}_{1})\leq\cdots\leq\dim(\mathcal{A}_{m})$ (note that one could use another refinement factor). For simplicity, we assume that $\mathcal{A}_{i}$ are subspaces of $\mathcal{A}$ so that the injection operator $\mathcal{A}_{i}\to\mathcal{A}$ is trivial. Let $X_{\theta}:\mathcal{A}\to\mathbb{R}$ be the operator defined as $X_{\theta}(a)=\|\mathcal{G}^{\dagger}(a)-\mathcal{G}_{\theta}(a)\|_{\mathcal{U}}^{2}$. We define the Galerkin projection operator $\mathcal{P}_{i}:\mathcal{A}\to\mathcal{A}_{i}$ and $X_{\theta}^{i}=X_{\theta}\circ\mathcal{P}_{i}$. Note that the neural operator $\mathcal{G}_{\theta}$ is discretization invariant following the framework introduced in [25], and that computing $X_{\theta}^{1}$ (and its gradient with respect to $\theta$) is inherently cheaper than computing $X_{\theta}^{m}$. Hence, the projection step to the latent dimension of the neural operator is computationally cheaper. Note that this applies for any neural operator architecture written in the encoder-processor-decoder framework [4, 23]. FNOs [32] could also be written in this framework by considering a spectral discretization of the function space, and the Fourier transform as the encoder.
Multi-Level Monte Carlo gradient estimation
The goal of the MLMC [15] gradient estimation is to approximate the expectation of the gradient of the loss function with respect to the neural network parameters $\theta$ using a telescopic sum of differences across mesh levels. The main idea is to use a coarse mesh to capture most of the variance in the gradient, and then use finer meshes to correct for the remaining variance. This allows us to reduce the number of fine-mesh samples needed, thus speeding up the training process. We express the fine-mesh gradient expectation, approximating $\mathbb{E}_{a\sim\mu}[\nabla_{\theta}X_{\theta}(a)]$, as a telescopic sum of differences across mesh levels and decompose the expectation of the gradient in Eq. 2 as follows:
@@
$\approx\underbrace{\frac{1}{N_{1}}\sum_{j=1}^{N_{1}}[\nabla_{\theta}X_{\theta}^{1}(a_{1,j})]}_{Y_{1}}+\sum_{i=2}^{m}\underbrace{\left(\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}[\nabla_{\theta}X_{\theta}^{i}(a_{i,j})-\nabla_{\theta}X_{\theta}^{i-1}(a_{i,j})]\right)}_{Y_{i}},$
@@
where $\{N_{i}\}_{1\leq i\leq m}$ denotes the sample sizes at each level. Note that we do not consider discretization of the output space $\mathcal{U}$ for simplicity, but a similar argument would also work, see, e.g., [4].
Optimal allocation of samples
Following [15, Sec. 1.3], the sequence $Y_{1}+\ldots+Y_{m}$ in the second line of Eq. 3 provides an unbiased estimator for $\mathbb{E}_{a\sim\mu}[\nabla_{\theta}X_{\theta}(a)]$, which approximates it with increasing accuracy, but also increasing cost, as we require more samples at the finest level of the mesh hierarchy. If we assume that each term inside the $Y_{i}$ can be evaluated at a computational cost of $C_{i}$ (where $C_{1}\leq\ldots\leq C_{m}$ as the mesh becomes finer), then the total cost of the estimator is
$C=\sum_{i=1}^{m}N_{i}C_{i}.$
We define the mean square error (MSE) of the estimator as
$\textup{MSE}=\mathbb{E}_{a_{i,j}\sim\mu}\left[\left(\sum_{i=1}^{m}Y_{i}-\mathbb{E}_{a\sim\mu}[\nabla_{\theta}X_{\theta}(a)]\right)^{2}\right],$
and aim to find an optimal allocation of samples $N_{1},\ldots,N_{m}$ that minimizes the total computational cost, while achieving a mean squared error of $\epsilon^{2}$, for some $\epsilon>0$. The following proposition relies on a standard MLMC result [14, 13, 15] and provides a mean squared error bound for the MLMC estimator in Eq. 3.
Proposition 2.2 (Sample allocation for gradient estimation).
@@
Let $\epsilon>0$ and suppose that the cost $C_{i}$ of each sample in Eq. 3 scales linearly with respect to the number of degrees of freedom in the finite element space as $C_{i}\leq c2^{di}$. Under 2.1, there exist a number of levels $m>0$ and an optimal allocation of samples $N_{i}\propto 2^{-(2k+d)i/2}$ per level such that the MLMC estimator to $\mathbb{E}_{a\sim\mu}[\nabla_{\theta}X_{\theta}(a)]$ given by Eq. 3 has a mean squared error bounded by $\epsilon^{2}$.
Proof.
Let $1\leq i\leq m$ be the level of the mesh. For a given $a\sim\mu$, we are interested in bounding the difference between the gradients on the coarse and fine meshes in Eq. 3. We remark that $X_{\theta}^{i}(a)=X_{\theta}(\mathcal{P}_{i}a)$, and for $a\sim\mu$ denote $H_{\theta}(a)\coloneqq\mathcal{G}^{\dagger}(a)-\mathcal{G}_{\theta}(a)$. Since $\mathcal{G}^{\dagger}$ and $\mathcal{G}_{\theta}$ are Lipschitz continuous, the application $H_{\theta}:\mathcal{A}\to\mathcal{U}$ is Lipschitz continuous with Lipschitz constant $\mathrm{Lip}(H_{\theta})=\mathrm{Lip}(\mathcal{G}^{\dagger})+\mathrm{Lip}(\mathcal{G}_{\theta})$. Let $1\leq s\leq p$ and $a\sim\mu$, then
$|\nabla_{\theta_{s}}X_{\theta}(\mathcal{P}_{i}a)-\nabla_{\theta_{s}}X_{\theta}(a)|$ $=2|\langle H_{\theta}(\mathcal{P}_{i}a),\nabla_{\theta_{s}}\mathcal{G}_{\theta}(\mathcal{P}_{i}a)\rangle_{\mathcal{U}}-\langle H_{\theta}(a),\nabla_{\theta_{s}}\mathcal{G}_{\theta}(a)\rangle_{\mathcal{U}}|$
$\leq 2|\langle H_{\theta}(a),\nabla_{\theta_{s}}\mathcal{G}_{\theta}(a)-\nabla_{\theta_{s}}\mathcal{G}_{\theta}(\mathcal{P}_{i}a)\rangle_{\mathcal{U}}|$
$\quad+2|\langle H_{\theta}(a)-H_{\theta}(\mathcal{P}_{i}a),\nabla_{\theta_{s}}\mathcal{G}_{\theta}(\mathcal{P}_{i}a)\rangle_{\mathcal{U}}|$
$\leq 2(L_{\theta}\|H_{\theta}(a)\|_{\mathcal{U}}+\mathrm{Lip}(H_{\theta})\|\nabla_{\theta_{s}}\mathcal{G}_{\theta}(\mathcal{P}_{i}a)\|_{\mathcal{U}})\|a-\mathcal{P}_{i}a\|_{\mathcal{A}}$
$\leq C_{\theta}\|a-\mathcal{P}_{i}a\|_{\mathcal{A}},$
where we used the Cauchy-Schwarz inequality to obtain the second inequality, and introduced the constant $C_{\theta}>0$ to bound the first term in the second inequality uniformly over $a$ because $\mu$ is compactly supported on $H^{k}(\Omega_{\mathcal{A}})$ and the operators are continuous. Here, we require that $\nabla_{\theta}\mathcal{G}_{\theta}$, $\mathcal{G}^{\dagger}$, and $\mathcal{G}_{\theta}$ are Lipschitz continuous, and use the fact that $a$ belongs to a compact domain and $H_{\theta}$, $\nabla_{\theta}\mathcal{G}_{\theta}$ are continuous to bound the last two terms uniformly over $a\sim\mu$. Finally, we employ a standard finite element error estimate [6, Eq. 4.4.28] to bound the difference between the projection $\mathcal{P}_{i}a$ and the original function $a$ as
Using the compactness of the measure $\mu$, we conclude that there exists a constant $M>0$ independent of $i$ such that
$\|\nabla_{\theta}X_{\theta}^{i}(a)-\nabla_{\theta}X_{\theta}(a)\|_{2}\leq M2^{-ki}.$ (4)
Moreover, by triangular inequality, we have
$\|\nabla_{\theta}X_{\theta}^{i}(a)-\nabla_{\theta}X_{\theta}^{i-1}(a)\|_{2}\leq\|\nabla_{\theta}X_{\theta}^{i}(a)-\nabla_{\theta}X_{\theta}(a)\|_{2}+\|\nabla_{\theta}X_{\theta}^{i-1}(a)-\nabla_{\theta}X_{\theta}(a)\|_{2}\leq(1+2^{k})M2^{-ki}.$ (5)
Combining Eqs. 4 and 5, we obtain the following bound for the expectation of the approximation of the gradient, and the variance of the multi-level estimator:
$\mathbb{E}_{a\sim\mu}[\|\nabla_{\theta}X_{\theta}^{i}(a)-\nabla_{\theta}X_{\theta}(a)\|_{2}]\leq M2^{-ki},\quad\textrm{Var}_{a\sim\mu}[\|\nabla_{\theta}X_{\theta}^{i}(a)-\nabla_{\theta}X_{\theta}^{i-1}(a)\|_{2}]\leq\frac{(1+2^{k})^{2}M^{2}}{4}2^{-2ki},$
where we used Popoviciu’s inequality to bound the variance of the bounded random variable $\|\nabla_{\theta}X_{\theta}^{i}(a)-\nabla_{\theta}X_{\theta}^{i-1}(a)\|_{2}$. Moreover, the assumption on the growth of the computational cost with respect to the number of levels implies that $C_{i}\leq c2^{di}$. Then, following [15, Thm. 2.1], for any $\epsilon>0$, there exists a number of level $m$, and optimal allocation of samples $N_{i}\propto 2^{-di/2}$ per level such that the MLMC estimator given by Eq. 3 has a mean square error bounded by $\epsilon^{2}$. ∎
@@
Following Proposition 2.2, we find that one can reduce the number of fine-mesh samples to accurately estimate the gradient of the loss function by adapting the number of samples at each level depending on the mesh resolution. Note that the number of samples per resolution is related to the sample cost assumption, which depends on the type of neural operator architecture. The analysis can also be generalized to non-geometric hierarchy of meshes [16].
@@
(Multi-Level Quasi-Monte Carlo) The computational cost of MLMC can be reduced by employing a Quasi-Monte Carlo (QMC) approach, where a nested set of samples across mesh resolutions is chosen instead of independent realizations at each level [26, 27]. This increases the variance of the estimator but results in a lower cost complexity [15].